OpenStudy (anonymous):

Find the distance D from the point A(0,-8,8) to the plane 3x + 7y - 2z + 5 = 0

OpenStudy (anonymous):

i did S(0, -8, 8) P would then be P(0, -5/7, 0).... just because it's an easy point. the y intercept \[\rightarrow PS\] = -51/7j + 8k magnitude of 3x + 7y-2z = -5 would be \[\sqrt{60}\] then just take dot product of ps and <3/\[\sqrt{60}\], 7/\[\sqrt{60}\], -2\[\sqrt{60}\] would be 0 - 15/\[\sqrt{60}\] - 67/\[\sqrt{60}\]

OpenStudy (anonymous):

er -15/sqr 60 - 16/sqrt 60 = -67/sqrt 60... absolute value so it would be 67 / sqrt 60